OpenStudy (anonymous):

Help !? Find the value of k for which the equation has a real double root: 3x2 - 6x - k = 0

OpenStudy (anonymous):

Are you familiar with the quadratic equation?

OpenStudy (anonymous):

yes but how do i use that if i dont know the value of k?

OpenStudy (anonymous):

k is your c.

OpenStudy (anonymous):

\[ax^2 + bx + c = 0 \implies x = {-(b) \pm \sqrt{(b)^2 - 4(a)(c)} \over 2(a)}\] So what is a, b, and c?

OpenStudy (anonymous):

From your problem I mean?

OpenStudy (anonymous):

A is3

OpenStudy (anonymous):

B is 6

OpenStudy (anonymous):

is the answer 33

OpenStudy (anonymous):

a = 3 b = -6 c = -k \[x = {-(-6) \pm \sqrt{(-6)^2 - 4(3)(-k)}\over 2(3)}\] So when will you get 2 values for x? When will you get 1?

OpenStudy (anonymous):

I'll give you a hint. What happens when the part inside the square root equals 0?

OpenStudy (anonymous):

doesnt it cancle out ? is it -3

OpenStudy (anonymous):

K= 2/3

OpenStudy (anonymous):

It would be x = 6/6 = 1. So that would have only one solution. So what does k need to be to make the part inside the square root 0?

OpenStudy (anonymous):

thanks i got it now :)